# Motion Manifold Flow Primitives for Language-Guided Trajectory Generation

### v1

@@ 7 Conclusion
Figure 10: Real robot waving motion generation by MMFP.
The Motion Manifold Flow Primitive (MMFP) framework presented in this paper combines motion manifolds and latent space flows, and allows for the construction of models that can learn complex text-conditional distributions from a small number of trajectory data, each labeled with multiple text annotations. Proper regularization has been designed to ensure the smoothness of the generated trajectories and robustness to unstructured text variations. Our experimental results demonstrate accurate motion generation even for unseen text inputs, which contrasts with existing manifold-based methods such as TCVAE and MMP, and trajectory space diffusion and flow-based models.

### v2

@@ 6 Conclusion
The Motion Manifold Flow Primitive (MMFP) framework combines motion manifolds and latent flows, enabling models to learn complex text-conditional distributions with dramatic changes in the number of modalities and support volumes based on text inputs, all while relying on a small number of trajectory data, each labeled with multiple text annotations. Proper regularization has been designed to ensure the smoothness of the generated trajectories and robustness to unstructured text variations. The performance of the MMFP framework has been validated through extensive experiments involving the generation of SE(3) and 7-DoF joint trajectories. Our results demonstrate accurate motion generation even for unseen text inputs generated by the Large Language Model, ChatGPT. This contrasts with existing manifold-based methods, such as TCVAE and MMP, which struggle with text inputs containing detailed descriptions, and trajectory space diffusion and flow-based models, which tend to produce jerky trajectories.
We believe the MMFP framework can be enhanced in several ways. The current implementation only uses text as input to the model. To develop more versatile movement primitives, incorporating visual inputs as an additional conditional vector will be crucial. For example, in a pouring task, the trajectory should depend on the type of bottle, whether it’s a kettle, a can, etc., with the bottle’s geometry provided as a depth image or point cloud. Developing models conditioned on both text and visual inputs may require more sophisticated neural network architectures to learn generalizable primitives from small datasets, which presents a challenging yet important problem.
Adopting a more structured text embedding space could enhance text generalization performance compared to the current vector space approach. Poincaré embeddings, known for effectively learning hierarchical representations (Nickel and Kiela, 2017), have shown promise in this regard. Given the hierarchical nature of text descriptions – such as the level 1, 2, and 3 texts discussed in our paper – hyperbolic embeddings can be particularly well-suited for capturing these relationships. Incorporating this type of structured geometry into our framework offers an intriguing direction for future research.
Our current framework relies on fixed-length discrete-time trajectory data, and thus demonstration trajectories must be pre-processed to have the same length. This requirement could potentially be relaxed by using sequence models, such as transformers or recurrent neural networks, which can handle variable-length data, for the decoder model, whereas the current implementation uses fully connected neural networks. Alternatively, as recently presented in (Lee, 2024), using continuous-time representations with parametric curve models is another promising direction.
{acks}
Yonghyeon Lee was the beneficiary of an individual grant from CAINS supported by a KIAS Individual Grant (AP092701) via the Center for AI and Natural Sciences at Korea Institute for Advanced Study. Byeongho Lee, Seungyeon Kim, and Frank C. Park were supported in part by IITP-MSIT grant RS-2021-II212068 (SNU AI Innovation Hub), IITP-MSIT grant RS-2022-II220480 (Training and Inference Methods for Goal Oriented AI Agents), IITP-MSIT grant RS-2024-00436680 (Collaborative Research Projects with Microsoft Research), KIAT grant P0020536 (HRD Program for Industrial Innovation), ATC+ MOTIE Technology Innovation Program grant 20008547, SRRC NRF grant RS-2023-00208052, SNU-AIIS, SNU-IPAI, SNU-IAMD, SNU BK21+ Program in Mechanical Engineering, and SNU Institute for Engineering Research.